OpenStudy (anonymous):

simplify: 1+sin2x - cos2x / 1+sin2x + cos2x

OpenStudy (anonymous):

Is it squared or "2x"?

OpenStudy (anonymous):

just 2x the way it's written?

OpenStudy (mertsj):

\[\frac{1+\sin 2x-\cos 2x}{1+\sin 2x+\cos 2x}=\frac{1+2\sin x \cos x+(1-2\sin ^2x)}{1+2\sin x \cos x+(2\cos ^2x-1)}\]

OpenStudy (mertsj):

\[\frac{2\sin ^2x+2\sin x \cos x}{2\cos ^2x+2\sin x \cos x}=\frac{2\sin x(\sin x+\cos x)}{2\cos x(\sin x+\cos x)}\]

OpenStudy (mertsj):

\[\tan x\]

OpenStudy (mertsj):

And i would point out that in the very first step, there is a sign error in the numerator. It should be 1+2sinxcosx-(1-2sin^2x)